Use the following price and quantity information is this table to answer the following questions. Just list the mathematical results to each question in the space below in SNoodle and show your work on your scratch paper 10 Price 40 Using 2005 as the base year, calculate the Real GDP in 2010 and 2011.
Answer #1

Real GDP use the base year prices to compute the final value of goods and services in different years. In this question the prices are fixed at 2005 prices. GDP is the sum of product of price and quantity of different goods.

Real GDP 2010 = 45*250 + 2200*2 + 310*50 = 31150

Real GDP 2011 = 50*250 + 2300*2 + 350*50 = 34600